Complete the following table by calculating the missing entries. In each case indicate whether the solution is acidic or basic.




pH pOH [H+] [OH -] acidic or basic?
12.40 M M

acidic

basic     

12.67 M M

acidic

basic     

2.40 x 10-13 M M

acidic

basic     

M 5.40 x 10-13 M

acidic

basic     
Answer #1

PR DH (1) PH = 12.40 POH= luopH = 1600 12.40 slob [at] = le pH = 10/2600 = 3.98% 1013 Sor] =10-POH = 10l6 = 0.025m . pH is gr